A 6.0-cm-diameter, 11-cm-long cylinder contains 100 mg of oxygen (O2) at a pressure less than 1 atm. The cap on one end of the cylinder is held in place only by the pressure of the air. One day when the atmospheric pressure is 100 kPa, it takes a 177 N force to pull the cap off.

What is the temperature of the gas?
Answer #1

Applied Force +Internal Pressure*Area =Atmospheric Pressure*Area

177N+P_{i}*pi r^{2}=P_{atm}*pi r^{2}

r = 0.06/2 0.03m

177N + P * π(0.03m)2-10°Pa * π(0.03m) rn rn

177N + P, * π(0.03m)--282.74 N

P, * π(0.03m)2-282. 74:V-177N

P, 105.74N/(π(0.03m)2)

P 3.74 * 10 Pa

=====================

No of moles,n 0.1g/(32g/mol)

n-3.125 103mol

--------------------

Use Formula .a1

where V is Volume =pi r^{2}h=pi (0.03m)^{2}(0.11m)

T=PV/nR

T=3.74*10^{4}Pa*pi (0.03m)^{2}(0.11m)/(3.125*10^{-3}mol*8.314J/K·mol)

T = 3.74 * 104 * π0 . 0320.11/(3.125 * 10-3 * 8.314)

ANSWER: T=447.71K

======================